The potential difference across a resting neuron in the human body is about 100.0 mV and carries a current of about 0.580 mA. How much power does the neuron release?

µW?
Answer #1

Here ,

Potential across neuron , v = 100 mV

current , I = 0.580 mA

I = 5.8 *10^-4 A

power released by neutron = V* I

power released by neutron = 100 *10^-3 * 5.8 *10^-4

power released by neutron = 5.8 *10^-5 W

power released by neutron = 58 uW

the power released by neutron is 58 uW